# Methodology for constraining ultralight vector bosons with gravitational wave searches targeting merger remnant black holes

### v1

@@ IV Weakly coupled vector bosons
The standard assumption in interpreting search results for signatures of BH superradiance and the presence of vector boson clouds is that the underlying massive vector field only acts gravitationally and has no couplings to the Standard Model or additional interactions. (In particular, one assumes that the vector mass arises through the Stueckelberg mechanism Lacki et al. (2009); Ruegg and Ruiz-Altaba (2004), as opposed to a Higgs mechanism.) The inference method and the simulation example discussed in earlier sections by default make this standard assumption. In this section, we outline how this assumption can be partially lifted, and how the constraints obtained from the approach described above can be mapped to the regime where additional interactions exist, but are not strong enough to nullify such searches. Specifically, we focus on a kinetically mixed dark photon (Sec. IV.1) and a dark Higgs-Abelian sector (Sec. IV.2) as two well-motivated examples of vector interactions. In Sec. IV.3, we forecast the parameter space accessible to next-generation gravitational detectors such as the Einstein Telescope and Cosmic Explorer.
@@ V Conclusion
For the synthetic system studied in Sec. III.2, we could exclude at best a fraction of one order of magnitude in the full vector boson mass range, assuming it was a real event and the search did not yield a detection. In reality, most CBC remnant BHs to date are still beyond reach for vector searches (or marginally reachable but not promising enough to derive constraints with high confidence). In future observing runs with further improved detectors, and in the era of next-generation detectors, we can expect a large population of CBC detections with much higher SNRs Abbott et al. (2023, 2020), which will enable sensitive searches for vector boson clouds hosted by the remnants. We also mention that recently an improved model for the frequency evolution of the GW signals from vector clouds has been obtained in Ref. May et al. (2024). Though the slightly increased estimates for the frequency time derivative obtained in this improved model (by $\sim 10\%$ for the typical systems considered here) will not affect the conclusions of this work, a more accurate frequency evolution model could not only be used to more optimally tune the configuration of the search strategy used here, but might also be leveraged to design more sensitive searches that track the signal coherently over longer timescales. By targeting remnant BHs with a variety of masses, we can correspondingly probe a wide boson mass range. Thus, this work merely lays the groundwork for the broader constraints we will be able to derive with future GW observations in a higher SNR regime.
@@ Appendix A Search configurations
For the search and follow-up study described in Secs. IIIII, we must choose a set of configurations that is both computationally feasible and provides sufficient coverage of the full parameter space. Because we use a flexible HMM-based search method and only track the signal frequency evolution over time, we find that a single configuration (with parameters $t_{\rm start}$, $T_{\rm coh}$, and $T_{\rm obs}$) is able to recover signals from systems with a range of parameters. An example is given in Fig. 5, where the detection probability is shown as the colored contours for synthetic injections generated by systems with a grid of parameters $M$ and $\alpha$ (all other parameters fixed to the medians shown in Table 1). The search configuration, optimal for a system with parameters marked by “$+$” in the figure, remains fixed across all injections: $t_{\rm start}-t_{0}=3.62$ hr, where $t_{0}$ is the time at which the BH was born, $T_{\rm coh}=7.2$ min, and $T_{\rm obs}=9.72$ hr. This single configuration is able to recover injected signals across the majority of the parameter space, even if, in most cases, it is not the optimal configuration. [^6] The optimal configuration refers to $t_{\rm start}=t_{\rm sat}$ (the saturation time of the superradiant cloud growth), $T_{\rm coh}\approx(2\dot{f}^{\rm max}_{\rm GW})^{-1/2}$ [see Eq. (1)], and $T_{\rm obs}\sim\tau_{\rm GW}$ (the GW emission timescale) Jones et al. (2023). As demonstrated in Sec. III, such a synthetic system is placed just within reach for us to derive meaningful constraints on $m_{V}$. The SNRs of the simulated signals are not so high that any mismatched search configuration would have recovered them. Hence, in practice, for any target system that our search is sensitive to, it is generally the case that a single search configuration can cover at least a range of the relevant parameter space.
@@
Nevertheless, the parameter space we consider in a real vector boson search is usually larger than that shown in the above example. Given the uncertainties associated with the remnant BH properties and the unknown boson mass range (or equivalently, the range of $\alpha$), it is beneficial to use multiple search configurations to provide better coverage and improve the overall search sensitivity. Still, we need to be conservative in selecting the configurations. Here we describe the procedure of selecting a set of search configurations for a given target, using the synthetic GW170814-like system discussed in Sec. III as an example. It is worth noting, however, that these guidelines should be generally applicable to the majority of merger remnants one might target in a vector boson search—that is, there is nothing special about the GW170814-like system we consider here.
@@ B.1.2 Critical kinetic mixing
That is, all signals with $\beta\lesssim\beta_{1}$ are detectable if $T_{\rm obs}$ is chosen based on the $\varepsilon=0$ expectations for the above fiducial BH remnant. In principle, a larger injection study would be required to confirm this more broadly. However, as we will see below, this is the less conservative condition on the kinetic mixing strength, so we refrain from investigating this condition across a larger set of search configurations.
@@ B.2.1 Modified frequency evolution
To obtain a more accurate estimate, albeit still with large uncertainties as discussed below, we perform numerical time-domain simulations of the Higgs-Abelian field equations on a Kerr BH background from appropriately chosen initial data as in Ref. East (2022). Here we focus on scenarios deep in the weakly nonlinear regime, i.e., at low field amplitudes. First, these simulations confirm that $\dot{E}_{\rm rad}$ is enhanced by higher-order nonlinearities as $A^{\prime 2}$ approaches $A^{\prime 2}_{c}$ before strings are formed. Second, isolating the leading quartic contribution to $\dot{E}_{\rm rad}$ (i.e., those that scale as $\sim M_{c}^{3}$), we find a steeper than $\sim\alpha^{6}$ scaling of the flux in the $\alpha\in(0.3,0.5)$ regime. This suggests that extrapolating these results to $\alpha\approx 0.1$ (relevant for this work) and using a $\sim\alpha^{6}$ scaling overestimates the flux in that regime. Lastly, using our evolution at $\alpha=0.3$, we can place the upper limit on the flux ratio, $\dot{E}_{\rm rad}A^{\prime 4}_{c}/\dot{E}_{\rm BH}^{3}\lesssim 6$, where $\dot{E}_{\rm BH}$ is the bosonic flux through the BH horizon (defined to be positive if energy is extracted from the BH, and which can be independently computed using SuperRad). Assuming this upper limit can be extrapolated as $\propto\alpha^{6}$, we obtain $\dot{E}_{\rm rad}\approx 3\times 10^{-9}\alpha^{6}A^{\prime-4}_{c}(M_{c}/M)^{3}$. This suggests that $\dot{E}_{\rm rad}/\dot{E}_{\rm BH}<1$ for $A^{\prime 2}\approx A^{\prime 2}_{c}$ and $\alpha\sim 0.1$; hence, the superradiance instability saturates gravitationally. Due to the uncertainties in our numerical time-domain evolution, the large extrapolation from the relativistic to the non-relativistic regimes, and the suggested steeper scaling with $\alpha$, this estimate of the flux is to be understood as an (at best) order-of-magnitude estimate. For the reasons above, this is likely a conservative estimate for $\alpha\lesssim 0.3$, i.e., it overestimates the total flux in the regime of parameter space relevant to this work and, hence, the impact on the GW frequency evolution.

### v2

@@ IV Weakly coupled vector bosons
The standard assumption in interpreting search results for signatures of BH superradiance and the presence of vector boson clouds is that the underlying massive vector field only acts gravitationally and has no couplings to the Standard Model or additional interactions. (In particular, one assumes that the vector mass arises through the Stueckelberg mechanism Lacki et al. (2009); Ruegg and Ruiz-Altaba (2004), as opposed to a Higgs mechanism.) The inference method and the simulation example discussed in earlier sections by default make this standard assumption. In this section, we outline how this assumption can be partially lifted, and how the constraints obtained from the approach described above can be mapped to the regime where additional interactions exist, but are not strong enough to nullify such searches. Specifically, we focus on a kinetically mixed dark photon and a dark Higgs-Abelian sector as two well-motivated examples of vector interactions.
There are several ways that bosonic interactions can affect the evolution, and hence the GW signal, of a superradiant cloud:
  1. i.
    The interactions can change how the energy, and hence oscillation frequency, of the boson cloud depends on the number of bosons in the cloud. This will change the frequency of the GW signal.
  2. ii.
    The interactions can cause the cloud to saturate at a lower total energy than it would if the superradiant instability were to saturate gravitationally (by spinning down the BH). This reduces the maximum amplitude of the resulting GW signal. This can happen due to the fact that bosonic interactions introduce additional dissipation channels, e.g., due to bosonic radiation, that eventually cancel out the superradiant extraction of energy through the BH.
  3. iii.
    Even if the extra dissipation channels due to the bosonic interactions are not strong enough to change the saturation of the superradiant instability, they can still affect how quickly the cloud dissipates away after saturation. If the bosonic dissipation is stronger or comparable to gravitational radiation, it will affect how long the GW signal lasts and how quickly the frequency evolves.
For the kinetically mixed dark photon (Sec. IV.1), we primarily consider when interactions with an electron-positron plasma cause the cloud to dissipate faster (effect iii). For the Higgs-Abelian model (Sec. IV.2), all three of these effects are relevant to the GW signal. We consider how interactions mediated by the Higgs-like field can change the self-energy of the cloud and lead to bosonic radiation, as well as how, in some parts of the parameter space, the growth of the cloud will be halted due to a stringy bosenova.
In Sec. IV.3, we forecast the parameter space accessible to next-generation GW detectors such as the Einstein Telescope and Cosmic Explorer.
@@ V Conclusion
For the synthetic system studied in Sec. III.2, we could exclude at best a fraction of one order of magnitude in the full vector boson mass range, assuming it was a real event and the search did not yield a detection. In reality, most CBC remnant BHs to date are still beyond reach for vector searches (or marginally reachable but not promising enough to derive constraints with high confidence). In future observing runs with further improved detectors, and in the era of next-generation detectors, we can expect a large population of CBC detections with much higher SNRs Abbott et al. (2023, 2020), which will enable sensitive searches for vector boson clouds hosted by the remnants. We also mention that recently an improved model for the frequency evolution of the GW signals from vector clouds has been obtained in Ref. May et al. (2024). Though the slightly increased estimates for the frequency time derivative obtained in this improved model (by $\sim 10\%$ for the typical systems considered here) will not affect the conclusions of this work, a more accurate frequency evolution model could not only be used to more optimally tune the configuration of the search strategy used here, but might also be leveraged to design more sensitive searches that track the signal coherently over longer timescales. By targeting remnant BHs with a variety of masses, one can correspondingly probe a wide boson mass range. Thus, this work merely lays the groundwork for the broader constraints that could be derived with future GW observations in a higher SNR regime.
@@ Appendix A Search configurations
For the search and follow-up study described in Secs. IIIII, one must choose a set of configurations that is both computationally feasible and provides sufficient coverage of the full parameter space. Because we use a flexible HMM-based search method and only track the signal frequency evolution over time, we find that a single configuration (with parameters $t_{\rm start}$, $T_{\rm coh}$, and $T_{\rm obs}$) is able to recover signals from systems with a range of parameters. Note, in this work we consider a signal detectable for a given configuration if, after running multiple noise realizations, the recovery rate is $\geq 90\%$, i.e., the false dismissal probability is $\leq 10\%$. An example is given in Fig. 5, where the detection probability is shown as the colored contours for synthetic injections generated by systems with a grid of parameters $M$ and $\alpha$ (all other parameters fixed to the medians shown in Table 1). The search configuration, optimal for a system with parameters marked by “$+$” in the figure, remains fixed across all injections: $t_{\rm start}-t_{0}=3.62$ hr, where $t_{0}$ is the time at which the BH was born, $T_{\rm coh}=7.2$ min, and $T_{\rm obs}=9.72$ hr. This single configuration is able to recover injected signals across the majority of the parameter space (with a 1% false alarm probability), even if, in most cases, it is not the optimal configuration. [^6] The optimal configuration refers to $t_{\rm start}=t_{\rm sat}$ (the saturation time of the superradiant cloud growth), $T_{\rm coh}\approx(2\dot{f}^{\rm max}_{\rm GW})^{-1/2}$ [see Eq. (1)], and $T_{\rm obs}\sim\tau_{\rm GW}$ (the GW emission timescale) Jones et al. (2023). As demonstrated in Sec. III, such a synthetic system is placed just within reach for us to derive meaningful constraints on $m_{V}$. The SNRs of the simulated signals are not so high that any mismatched search configuration would have recovered them. Hence, in practice, for any target system that our search is sensitive to, it is generally the case that a single search configuration can cover at least a range of the relevant parameter space.
@@
Nevertheless, the parameter space considered in a real vector boson search is usually larger than that shown in the above example. Given the uncertainties associated with the remnant BH properties and the unknown boson mass range (or equivalently, the range of $\alpha$), it is beneficial to use multiple search configurations to provide better coverage and improve the overall search sensitivity. Still, one must be conservative in selecting the configurations. Here we describe the procedure of selecting a set of search configurations for a given target, using the synthetic GW170814-like system discussed in Sec. III as an example. It is worth noting, however, that these guidelines should be generally applicable to the majority of merger remnants one might target in a vector boson search—that is, there is nothing special about the GW170814-like system we consider here.
@@ B.1.2 Critical kinetic mixing
That is, all signals with $\beta\lesssim\beta_{1}$ are detectable if $T_{\rm obs}$ is chosen based on the $\varepsilon=0$ expectations for the above fiducial BH remnant. In principle, a larger injection study would be required to confirm this more broadly. However, as shown below, this is the less conservative condition on the kinetic mixing strength, so we refrain from investigating this condition across a larger set of search configurations.
@@ B.2.1 Modified frequency evolution
To obtain a more accurate estimate, albeit still with large uncertainties as discussed below, we perform numerical time-domain simulations of the Higgs-Abelian field equations on a Kerr BH background from appropriately chosen initial data as in Ref. East (2022). Here we focus on scenarios deep in the weakly nonlinear regime, i.e., at low field amplitudes. First, these simulations confirm that $\dot{E}_{\rm rad}$ is enhanced by higher-order nonlinearities as $A^{\prime 2}$ approaches $A^{\prime 2}_{c}$ before strings are formed. Second, isolating the leading quartic contribution to $\dot{E}_{\rm rad}$ (i.e., those that scale as $\sim M_{c}^{3}$), one finds a steeper than $\sim\alpha^{6}$ scaling of the flux in the $\alpha\in(0.3,0.5)$ regime. This suggests that extrapolating these results to $\alpha\approx 0.1$ (relevant for this work) and using a $\sim\alpha^{6}$ scaling overestimates the flux in that regime. Lastly, using our evolution at $\alpha=0.3$, we can place the upper limit on the flux ratio, $\dot{E}_{\rm rad}A^{\prime 4}_{c}/\dot{E}_{\rm BH}^{3}\lesssim 6$, where $\dot{E}_{\rm BH}$ is the bosonic flux through the BH horizon (defined to be positive if energy is extracted from the BH, and which can be independently computed using SuperRad). Assuming this upper limit can be extrapolated as $\propto\alpha^{6}$, we obtain $\dot{E}_{\rm rad}\approx 3\times 10^{-9}\alpha^{6}A^{\prime-4}_{c}(M_{c}/M)^{3}$. This suggests that $\dot{E}_{\rm rad}/\dot{E}_{\rm BH}<1$ for $A^{\prime 2}\approx A^{\prime 2}_{c}$ and $\alpha\sim 0.1$; hence, the superradiance instability saturates gravitationally. Due to the uncertainties in our numerical time-domain evolution, the large extrapolation from the relativistic to the non-relativistic regimes, and the suggested steeper scaling with $\alpha$, this estimate of the flux is to be understood as an (at best) order-of-magnitude estimate. For the reasons above, this is likely a conservative estimate for $\alpha\lesssim 0.3$, i.e., it overestimates the total flux in the regime of parameter space relevant to this work and, hence, the impact on the GW frequency evolution.